# MCMC Marginalisation Bias and $\Lambda$CDM tensions

### v2

@@ 2.2 Cosmic Chronometer (CC) Data
Here we work with OHD from the cosmic chronometer (CC) program Jimenez:2001gg . Concretely, we study 34 $H(z)$ constraints spanning the redshift range $0.07\leq z\leq 1.965$ Stern:2009ep ; Moresco:2012jh ; Zhang:2012mp ; Moresco:2016mzx ; Ratsimbazafy:2017vga ; Borghi:2021rft ; Jiao:2022aep ; Tomasetti:2023kek . We illustrate the data in Fig. 1, where it is consistent with Fig. 9 of Tomasetti:2023kek modulo the fact that we have an additional data point at $z=0.8$, which is not independent. See Table 1.1 of Moresco:2023zys for a discussion on the independence of the data points. While CC data may eventually be good enough to arbitrate on Hubble tension Moresco:2023zys , the data is not good enough on its own to do cosmology. To put this comment in context, we observe that the errors in Fig. 1 do not include systematic errors (see Moresco:2020fbm for an account of the systematics). As a result the constraints we get on cosmological parameters will be underestimated. Thus, from our perspective the data in Fig. 1 is simply some representative cosmological data in the OHD class.
@@
We impose a low redshift cut-off on the OHD $z_{\textrm{min}}$, removing all data points with redshifts $z_{i}<z_{\textrm{min}}$, and then extremising the $\chi^{2}$ likelihood,
@@
where $H_{i}:=H(z_{i})$ denotes OHD and $H_{\textrm{model}}(z)$ is the model (1) without the high redshift limit. The best fit $(H_{0},\Omega_{m})$ parameters correspond to the minimum of the $\chi^{2}$, while on the assumption of Gaussian errors, we estimate the errors from a Fisher matrix (appendix A). In parallel, we perform MCMC marginalisation through emcee Foreman-Mackey:2012any . More concretely, subject to the priors $H_{0}\in[0,200]$ and $\Omega_{m}\in[0,1]$, the latter restricting us to a physical regime, we record the peak and $68\%$ confidence interval from projected 1D $H_{0}$ and $\Omega_{m}$ posteriors.  [^7] We define the $68\%$ confidence interval by converting the 1D posterior into a histogram. Therein, each bin centered on a given $H_{0}$ or $\Omega_{m}$ value has a width $w_{i}$ and a count $c_{i}$. The total area of the discretised 1D posterior (histogram) is simply $A_{\textrm{tot}}=\sum_{i}w_{i}c_{i}$, where the sum ranges over all bins. Starting from the maximum count (posterior peak) in the histogram $c_{\textrm{max}}$, we identify the subset of bins with counts $c_{j}$ greater than a fixed value so that the normalised area is 68% of the total, i. e. $\sum_{j}w_{j}c_{j}/A_{\textrm{tot}}\sim 0.68$. The range of $H_{0}$ or $\Omega_{m}$ values covered by this subset of bins then defines the $68\%$ confidence interval. We have checked that our methodology agrees with GetDist Lewis:2019xzd and recovers the standard practice of quoting $16^{\textrm{th}}$, $50^{\textrm{th}}$ (peak) and $84^{\textrm{th}}$ percentiles for Gaussian posteriors. We stress that our Fisher matrix analysis assumes that 1D posteriors are Gaussian. As we shall see from our MCMC analysis, this is a good approximation when $z_{\textrm{min}}=0$, but as $z_{\textrm{min}}$ increases, 1D posteriors becomes more non-Gaussian due to a degeneracy giving rise to a banana-shaped contour in the 2D $(H_{0},\Omega_{m})$ posterior. The naive interpretation of this result is that the data is not good enough to constrain the model. As we shall argue, this interpretation is false. Nevertheless, by evaluating the MCMC chains on the $\chi^{2}$ likelihood (2), we confirm that the parameters corresponding to the minimum $\chi^{2}$ are tracking the best fit from the Fisher matrix analysis. This provides a non-trivial consistency check between the Fisher matrix analysis based on optimisation (gradient descent) and MCMC marginalisation, where the $\chi^{2}$ is simply evaluated on the MCMC chain.
@@
Table 1: Comparison between Fisher matrix and MCMC analysis for CC data with a low redshift cut-off $z_{\textrm{min}}$. We record the number of data points, the minimum of the $\chi^{2}$ (Fisher Matrix), the peak of the posteriors (MCMC), the $68\%$ ($1\sigma$) confidence intervals estimated from both the Fisher matrix and MCMC posteriors, and the minimum $\chi^{2}$ from the MCMC chain in brackets. MCMC marginalisation exhibits non-Gaussian $68\%$ confidence intervals, and for $z_{\textrm{min}}>1$, the minimum value of the $\chi^{2}$ from the MCMC chain falls outside of this interval. The $\chi^{2}$ minimum from the MCMC chain agrees with direct $\chi^{2}$ minimisation up to small numbers in line with expectations.
@@ 2.3 Features in CC Data
Turning our attention from the best fits to the MCMC posteriors, we see that the central values corresponding to the peaks of $H_{0}$ and $\Omega_{m}$ posteriors track the best fits with reasonably good agreement through to $z_{\textrm{min}}=0.6$. From $z_{\textrm{min}}=0.7$ onwards we note that the minimum of the $\chi^{2}$ appears near the boundary of the $H_{0}$ posterior $68\%$ confidence interval or beyond. Moreover, from the $\Omega_{m}$ posteriors, it is clear that there are long non-Gaussian tails in the direction of larger $\Omega_{m}$ values, which are in line with expectations when one fits the $\Lambda$CDM model to binned OHD at higher redshifts Colgain:2022tql . Interestingly, at $z_{\textrm{min}}=0.8$ and $z_{\textrm{min}}=1$ (see Fig. 2), the $H_{0}$ posterior is bimodal with a secondary peak at smaller $H_{0}$ values. Increasing $z_{\textrm{min}}$ further, this secondary peak, which is an artefact of a projection effect, dominates (see later Fig. 6) and explains the discrepancy between the peak of the $H_{0}$ posterior and the best fit for $z_{\textrm{min}}=1.2$ and $z_{\textrm{min}}=1.4$. Evidently, the $H_{0}$ posterior is being pulled to lower values by a degeneracy in the 2D posterior. For $z_{\textrm{min}}=1.5$ the $H_{0}$ peak and best fit agree well, but this is simply an accident due to a genuine low $H_{0}$ best fit.
@@
Evidently, given the non-Gaussian 1D posteriors, care is required when interpreting the significance of the trend towards a non-evolving (horizontal) $H(z)$ at higher redshifts in Fig. 1. We cannot use the errors from the Fisher matrix as we are clearly in a non-Gaussian regime, whereas MCMC inferences are impacted by projection effects to the extent that the minimum of the $\chi^{2}$ (confirmed from the MCMC chain) appears at the boundary of the $68\%$ confidence interval. For this reason, we resort to mock simulations. While this may seem a little redundant if we are going to employ profile distributions in section 3, there is motivation for this exercise. In Colgain:2022rxy the significance of a descending $H_{0}$/increasing $\Omega_{m}$ trend with effective redshift in OHD, Type Ia SN and QSOs was estimated to be a $\sim 3\sigma$ effect on the basis of combining $\sim 2\sigma$ effects in each of the independent data sets using Fisher’s method. Here, working with the same data throughout, we can directly compare the significance of a discrepancy estimated through mock simulations from the significance of a discrepancy estimated through profile distributions. In particular, we will address the question: how significant is a constant $H(z)$ with $z_{\textrm{min}}=1$ (8 data points) against the Planck consistent cosmology favoured by the full data set ($z_{\textrm{min}}=0$ entry in Table 1)? Note, the significance will be overestimated due to missing systematic uncertainties (see Moresco:2020fbm ), but we can still make comparison between mock simulations and profile distributions.
@@ 3 Profile Distributions
Having explained the mathematics behind the bias, which gives rise to a projection effect, in subsection 2.1, and having illustrated how it affects MCMC inferences in subsection 2.2 - the minimum of the $\chi^{2}$ risks falling outside of $68\%$ confidence intervals - we turn to profile distributions (PDs) Gomez-Valent:2022hkb , an extension of the profile likelihood, e. g. Trotta:2017wnx , to provide a complementary viewpoint. Consider two sets of parameters $\theta_{1}$ and $\theta_{2}$ and a normalised distribution $\mathcal{P}(\theta_{1},\theta_{2})$. The basic idea Gomez-Valent:2022hkb is to study the ratio
@@
The advantage of this approach is that $R(\theta_{1})$ can serve as a probability distribution function (up to an overall normalization), however we do not need to perform any integration, so $R(\theta_{1})$ is not prone to volume or projection effects. At this juncture, given the simplicity of our setup with only two parameters $(H_{0},\Omega_{m})$, we can be more explicit. Consider the probability distribution,
@@
We now apply the PD methodology to the non-Gaussian regime where MCMC marginalisation leads to biased results. To be concrete, we focus on the eight data points in the range $21<z<$ 2 where a non-evolving $H(z)$ trend is evident in the raw data in Fig. 1. Our goal here is to quantify the disagreement with the full data set, where one infers $H_{0}\sim 68.2$ km/s/Mpc and $\Omega_{m}\sim 0.32$. A similar exercise was performed in subsection 2.3 with mock simulations and the disagreement was estimated to be approximately $2\sigma$. Repeating the steps outlined above for the CC data with $z_{\textrm{min}}=1$ we find the distributions in Fig. 5. The first observation is that the distributions are non-Gaussian, but a comparison to the MCMC posteriors from the same data in blue in Fig. 2 reveals that there is no secondary $H_{0}$ peak at $H_{0}\sim 50$ km/s/Mpc. Thus, we confirm the secondary peak to be a projection effect. That being said, the primary $H_{0}$ peak from Fig. 2 has shifted to the dashed line corresponding to the minimum of the $\chi^{2}$, since the peak of the $R(H_{0})$ distribution and $\chi^{2}$ minimum agree by construction. Comparing the blue $\Omega_{m}$ distribution from Fig. 2 to the $R(\Omega_{m})$ distribution in Fig. 5, we see that the peak is close to $\Omega_{m}=0$ and that the tails continue to $\Omega_{m}=1$. In both plots we see that there is a non-zero probability of inferring $\Omega_{m}=1$. In some sense, this is not so surprising, the reason being that one is free to adopt generous priors for $H_{0}$, so that probability of large and small $H_{0}$ values is zero, but the priors on $\Omega_{m}$ in the flat $\Lambda$CDM model are restricted. For this reason, as a distribution spreads one invariably finds that distributions are impacted by the $\Omega_{m}$ priors. [^9] Note, this is a problem for the flat $\Lambda$CDM model. In particular, one may easily find that the peak of the $\Omega_{m}$ distribution is larger than $\Omega_{m}=1$, as is the case with Hubble Space Telescope SN with redshifts $z>1$ in the Pantheon+ sample Malekjani:2023dky .
@@
It is evident from Fig. 5 that any tension that exists is confined to the $H_{0}$ parameter. Moreover, since there may be only one binned value of $\Omega_{m}$ below the $R(\Omega_{m})$ peak, at the precision afforded to us by 200 bins, the $R(\Omega_{m})$ distribution in Fig. 5 is essentially one-sided and the $68\%$ confidence interval stretches beyond $\Omega_{m}\sim 0.32$, so there is no disagreement in the $\Omega_{m}$ parameter. Nevertheless, in the $H_{0}$ parameter we see that $H_{0}\sim 68.2$ km/s/Mpc, the value favoured by the full data set is at the boundary of the $95\%$ confidence interval, i.e. just under $2\sigma$ removed from the peak. The main point here is that, as is obvious from the raw data, current CC data with $z>1$ has a preference for a non-evolving Hubble parameter $H(z)$ with a large constant $H_{0}\sim 150$ km/s/Mpc. The disagreement is just under $2\sigma$, more accurately a $94\%$ ($1.9\sigma$) confidence level difference from $R(H_{0})$, and only $63\%$ ($0.9\sigma$) confidence level difference from $R(\Omega_{m})$. Although this may not be a serious discrepancy, essentially because of the poor data quality (8 data points), this disagreement supports the $\sim 2\sigma$ discrepancy seen in the mock simulations. It should be borne in mind that systematic uncertainties have been omitted and these will reduce this discrepancy once properly propagated. Given the agreement between the PD and mock simulation analysis, there is nothing to suggest that the three independent trends highlighted in Colgain:2022rxy across OHD, Type Ia SN and QSOs are not bona fide disagreements and that redshift evolution is present in the sample. The task remains to combine them at the level of a $\chi^{2}$ likelihood instead of combining them using Fisher’s method on the basis that they are independent probabilities. We leave this exercise for a forthcoming paper, but revisit the tension in OHD data in section 5. For completeness, in Table 2 we perform a reanalysis of CC data subsets with the PD approach and record the $68\%$ confidence intervals.

### v3

@@ 2.2 Cosmic Chronometer (CC) Data
Here we work with OHD from the cosmic chronometer (CC) program Jimenez:2001gg . Concretely, we study 34 $H(z)$ constraints spanning the redshift range $0.07\leq z\leq 1.965$ Stern:2009ep ; Moresco:2012jh ; Zhang:2012mp ; Moresco:2016mzx ; Ratsimbazafy:2017vga ; Borghi:2021rft ; Jiao:2022aep ; Tomasetti:2023kek . We illustrate the data in Fig. 1, where it is consistent with Fig. 9 of Tomasetti:2023kek modulo the fact that we have an additional data point at $z=0.8$, which is not independent. See Table 1.1 of Moresco:2023zys for a discussion on the independence of the data points. While CC data may eventually be good enough to arbitrate on Hubble tension Moresco:2023zys , the data is not good enough on its own to do cosmology. To put this comment in context, we observe that the errors in Fig. 1 do not include systematic errors (see Moresco:2020fbm for an account of the systematics). As a result the constraints we get on cosmological parameters will be underestimated. Thus, from our perspective the data in Fig. 1 is simply some representative cosmological data in the OHD class, providing a concrete setting where we can compare Bayesian and frequentist methods.
@@
We impose a low redshift cut-off on the OHD $z_{\textrm{min}}$, removing all data points with redshifts $z_{i}<z_{\textrm{min}}$, and then extremise the $\chi^{2}$ likelihood,
@@
where $H_{i}:=H(z_{i})$ denotes OHD and $H_{\textrm{model}}(z)$ is the model (1) without the high redshift limit. The best fit $(H_{0},\Omega_{m})$ parameters correspond to the minimum of the $\chi^{2}$, while on the assumption of Gaussian errors, we estimate the errors from a Fisher matrix (appendix A). In parallel, we perform MCMC marginalisation through emcee Foreman-Mackey:2012any . More concretely, subject to the priors $H_{0}\in[0,200]$ and $\Omega_{m}\in[0,1]$, the latter restricting us to a physical regime, we record the peak and $68\%$ credible interval from projected 1D $H_{0}$ and $\Omega_{m}$ posteriors.  [^7] We define the $68\%$ credible interval by converting the 1D posterior into a histogram. Therein, each bin centered on a given $H_{0}$ or $\Omega_{m}$ value has a width $w_{i}$ and a count $c_{i}$. The total area of the discretised 1D posterior (histogram) is simply $A_{\textrm{tot}}=\sum_{i}w_{i}c_{i}$, where the sum ranges over all bins. Starting from the maximum count (posterior peak) in the histogram $c_{\textrm{max}}$, we identify the subset of bins with counts $c_{j}$ greater than a fixed value so that the normalised area is 68% of the total, i. e. $\sum_{j}w_{j}c_{j}/A_{\textrm{tot}}\sim 0.68$. The range of $H_{0}$ or $\Omega_{m}$ values covered by this subset of bins then defines the $68\%$ credible interval. We have checked that our methodology agrees with GetDist Lewis:2019xzd and recovers the standard practice of quoting $16^{\textrm{th}}$, $50^{\textrm{th}}$ (peak) and $84^{\textrm{th}}$ percentiles for Gaussian posteriors. We stress that our Fisher matrix analysis assumes that 1D posteriors are Gaussian. As we shall see from our MCMC analysis, this is a good approximation when $z_{\textrm{min}}=0$, but as $z_{\textrm{min}}$ increases, 1D posteriors becomes more non-Gaussian due to a degeneracy giving rise to a banana-shaped contour in the 2D $(H_{0},\Omega_{m})$ posterior. The naïve interpretation of this result is that the data is not good enough to constrain the model. As we shall argue, this interpretation may only be true in the Bayesian framework. Nevertheless, by evaluating the MCMC chains on the $\chi^{2}$ likelihood (2), we confirm that the parameters corresponding to the minimum $\chi^{2}$ are tracking the best fit from the Fisher matrix analysis. This provides a non-trivial consistency check between the Fisher matrix analysis based on optimisation (gradient descent) and MCMC marginalisation, where the $\chi^{2}$ is simply evaluated on the MCMC chain.
@@
Table 1: Comparison between Fisher matrix and MCMC analysis for CC data with a low redshift cut-off $z_{\textrm{min}}$. We record the number of data points, the minimum of the $\chi^{2}$ (Fisher Matrix), the peak of the posteriors (MCMC), the $68\%$ ($1\sigma$) intervals estimated from both the Fisher matrix and MCMC posteriors, and the minimum $\chi^{2}$ from the MCMC chain in brackets. MCMC marginalisation exhibits non-Gaussian $68\%$ credible intervals, and for $z_{\textrm{min}}>1$, the minimum value of the $\chi^{2}$ from the MCMC chain falls outside of this interval. The $\chi^{2}$ minimum from the MCMC chain agrees with direct $\chi^{2}$ minimisation up to small numbers in line with expectations.
@@ 2.3 Features in CC Data
Turning our attention from the best fits to the MCMC posteriors, we see that the central values corresponding to the peaks of $H_{0}$ and $\Omega_{m}$ posteriors track the best fits with reasonably good agreement through to $z_{\textrm{min}}=0.6$. From $z_{\textrm{min}}=0.7$ onwards we note that the minimum of the $\chi^{2}$ appears near the boundary of the $H_{0}$ posterior $68\%$ credible interval or beyond. Moreover, from the $\Omega_{m}$ posteriors, it is clear that there are long non-Gaussian tails in the direction of larger $\Omega_{m}$ values, which are in line with expectations when one fits the $\Lambda$CDM model to binned OHD at higher redshifts Colgain:2022tql . Interestingly, at $z_{\textrm{min}}=0.8$ and $z_{\textrm{min}}=1$ (see Fig. 2), the $H_{0}$ posterior is bimodal with a secondary peak at smaller $H_{0}$ values. Increasing $z_{\textrm{min}}$ further, this secondary peak, which is an artefact of a projection effect, dominates (see later Fig. 7) and explains the discrepancy between the peak of the $H_{0}$ posterior and the best fit for $z_{\textrm{min}}=1.2$ and $z_{\textrm{min}}=1.4$. Evidently, the $H_{0}$ posterior is being pulled to lower values by a degeneracy in the 2D posterior. For $z_{\textrm{min}}=1.5$ the $H_{0}$ peak and best fit agree well, but this is simply an accident due to a genuine low $H_{0}$ best fit.
@@
Evidently, given the non-Gaussian 1D posteriors, care is required when interpreting the significance of the trend towards a non-evolving (horizontal) $H(z)$ at higher redshifts in Fig. 1. We cannot use the errors from the Fisher matrix as we are clearly in a non-Gaussian regime, whereas MCMC inferences are impacted by projection effects. For this reason, we resort to mock simulations. While this may seem a little redundant if we are going to employ profile distributions in section 3, there is motivation for this exercise. In Colgain:2022rxy the significance of a descending $H_{0}$/increasing $\Omega_{m}$ trend with effective redshift in OHD, Type Ia SNe and QSOs was estimated to be a $\sim 3\sigma$ effect on the basis of combining $\sim 2\sigma$ effects in each of the independent data sets using Fisher’s method. Here, working with the same data throughout, we can directly compare the significance of a discrepancy estimated through mock simulations from the significance of a discrepancy estimated through profile distributions. In particular, we will address the question: how significant is a constant $H(z)$ with $z_{\textrm{min}}=1$ (8 data points) against the Planck consistent cosmology favoured by the full data set ($z_{\textrm{min}}=0$ entry in Table 1)? Note, the significance will be overestimated due to missing systematic uncertainties (see Moresco:2020fbm ), but we can still make comparison between mock simulations and profile distributions.
@@ 3 Profile Distributions
Having explained the mathematics behind the bias, which gives rise to a projection effect, in subsection 2.1, and having illustrated how it affects MCMC inferences in subsection 2.2 - the minimum of the $\chi^{2}$ risks falling outside of $68\%$ credible intervals - we turn to profile distributions (PDs) Gomez-Valent:2022hkb , an extension of the profile likelihood, e. g. Trotta:2017wnx , to provide a complementary viewpoint. Consider two sets of parameters $\theta_{1}$ and $\theta_{2}$ and a normalised distribution $\mathcal{P}(\theta_{1},\theta_{2})$. The basic idea Gomez-Valent:2022hkb is to study the ratio
@@
The advantage of this approach is that $R(\theta_{1})$ can serve as a probability distribution function (up to an overall normalisation), however we do not need to perform any integration, so $R(\theta_{1})$ is not prone to volume or projection effects. It is worth noting that $R(\theta_{1})$ is a profile likelihood ratio in more common terminology Trotta:2017wnx . At this juncture, given the simplicity of our setup with only two parameters $(H_{0},\Omega_{m})$, we can be more explicit. Consider the probability distribution,
@@
As further remarks, we note that when profile likelihoods are Gaussian (see discussion in Colgain:2024clf ), as is visibly the case here in Fig. 4, it is valid to apply Wilks’ theorem Wilks:1938dza . Doing so, $68\%$ confidence intervals correspond to regions with $\Delta\chi^{2}\leq 1\Leftrightarrow R(\theta_{1})\gtrsim 0.6$. One can now draw a horizontal line across from $R(\theta_{1})\sim 0.6$, $\theta_{1}\in\{H_{0},\Omega_{m}\}$, in Fig. 4 and confirm that the line intersects $R(\theta_{1})$ at the dashed lines demarcating $68\%$ confidence intervals. The key point here is that despite the methodology in Gomez-Valent:2022hkb being new, the resulting confidence intervals agree with standard methods when distributions are Gaussian. It should be noted that Gomez-Valent:2022hkb bins the MCMC chain to construct PDs, which are essentially profile likelihood ratios, but this is acceptable, as explained in Trotta:2017wnx . As shown in Fig. 4 of Colgain:2024clf , for converged MCMC chains, the agreement between profile likelihoods constructed through optimisation and profile likelihoods constructed through binned MCMC chains is good. Here, in contrast to Colgain:2024clf , we run very long chains, well beyond the point at which they converge, and our model has fewer nuisance parameters, so the agreement is expected to be excellent.
@@
We now apply the PD methodology to the non-Gaussian regime where MCMC marginalisation leads to biased results and Wilks’ theorem Wilks:1938dza is no longer valid. To be concrete, we focus on the eight data points in the range $21<z<$ 2 where a non-evolving $H(z)$ trend is evident in the raw data in Fig. 1. Our goal here is to quantify the disagreement with the full data set, where one infers $H_{0}\sim 68.2$ km/s/Mpc and $\Omega_{m}\sim 0.32$. A similar exercise was performed in subsection 2.3 with mock simulations and the disagreement was estimated to be approximately $2\sigma$. Repeating the steps outlined above for the CC data with $z_{\textrm{min}}=1$ we find the distributions in Fig. 5. The first observation is that the distributions are non-Gaussian, but a comparison to the MCMC posteriors from the same data in blue in Fig. 2 reveals that there is no secondary $H_{0}$ peak at $H_{0}\sim 50$ km/s/Mpc. Thus, we confirm the secondary peak to be a projection effect. That being said, the primary $H_{0}$ peak from Fig. 2 has shifted to the dashed line corresponding to the minimum of the $\chi^{2}$, since the peak of the $R(H_{0})$ distribution and $\chi^{2}$ minimum agree by construction. Comparing the blue $\Omega_{m}$ distribution from Fig. 2 to the $R(\Omega_{m})$ distribution in Fig. 5, we see that the peak is close to $\Omega_{m}=0$ and that the tails continue to $\Omega_{m}=1$. In both plots we see that there is a non-zero probability of inferring $\Omega_{m}=1$. In some sense, this is not so surprising, the reason being that one is free to adopt generous priors for $H_{0}$, so that probability of large and small $H_{0}$ values is zero, but the priors on $\Omega_{m}$ in the flat $\Lambda$CDM model are restricted. For this reason, as a distribution spreads one invariably finds that distributions are impacted by the $\Omega_{m}$ priors. [^9] Note, this is a problem for the flat $\Lambda$CDM model. In particular, one may easily find that the peak of the $\Omega_{m}$ distribution is larger than $\Omega_{m}=1$, as is the case with Hubble Space Telescope SNe with redshifts $z>1$ in the Pantheon+ sample Malekjani:2023dky .
@@
It is evident from Fig. 5 that any tension that exists is confined to the $H_{0}$ parameter. Moreover, since there may be only one binned value of $\Omega_{m}$ below the $R(\Omega_{m})$ peak, at the precision afforded to us by 200 bins, the $R(\Omega_{m})$ distribution in Fig. 5 is essentially one-sided and the $68\%$ confidence interval stretches beyond $\Omega_{m}\sim 0.32$, so there is no disagreement in the $\Omega_{m}$ parameter. Nevertheless, in the $H_{0}$ parameter we see that $H_{0}\sim 68.2$ km/s/Mpc, the value favoured by the full data set is at the boundary of the $95\%$ confidence interval, i.e. just under $2\sigma$ removed from the peak. The main point here is that, as is obvious from the raw data, current CC data with $z>1$ has a preference for a non-evolving Hubble parameter $H(z)$ with a large constant $H_{0}\sim 150$ km/s/Mpc. The disagreement is just under $2\sigma$, more accurately a $94\%$ ($1.9\sigma$) confidence level difference from $R(H_{0})$, and only $63\%$ ($0.9\sigma$) confidence level difference from $R(\Omega_{m})$.
Note, since $R(\Omega_{m})$ is curtailed by the bounds on $\Omega_{m}$, inferences based on $\Omega_{m}$ need to be taken with a pinch of salt as they depend on our bounds. Moreover, from $R(H_{0})$ we note that the $1^{\textrm{st}}$ derivative fails to be continuous at $H_{0}\sim 40$ km/s/Mpc, which also appears to be an artefact of the $\Omega_{m}\leq 1$ bound. This kink happens beyond the $95\%$ confidence intervals, so even if corrected by relaxing the bounds on $\Omega_{m}$ it is not expected to greatly impact the results. We note that while it is not correct to apply Wilks’ theorem, the $R(H_{0})\sim 0.6$ horizontal line appears to intersect $R(H_{0})$ at the dashed lines denoting $68\%$ confidence intervals. This is not the case with $R(\Omega_{m})$, where the bounds are impacting results. We note that one can use the Feldman-Cousins prescription Feldman:1997qc to generalise methodology based on Wilks’ theorem when boundaries are present, but once again the methodology is only valid for Gaussian profiles, which is demonstrably not the case here. Ultimately, the methods in the literature rest upon Gaussian distributions being recovered in the large sample limit. In cosmology, whether one is in the large sample limit or not can only be judged once one has constructed the MCMC posterior or profile likelihood. Blindly applying theorems beyond their scope is neither good mathematics nor good physics.
Figure 6: $H_{0}$ constraints versus $z_{\textrm{min}}$ from Table 1 and Table 2. MCMC constraints follow Fisher matrix and profile constraints well at $z_{\textrm{min}}\lesssim 0.6$ and at $z_{\textrm{min}}=1.5$, but the credible intervals shift with respect to the confidence intervals to lower values at intermediary $z_{\textrm{min}}$. For presentation purposes we have shifted the constraints in $z_{\textrm{min}}$.
Although $\sim 2\sigma$ may not be a serious discrepancy, essentially because of the poor data quality (8 data points), this disagreement supports the $\sim 2\sigma$ discrepancy seen in the mock simulations. It should be borne in mind that systematic uncertainties have been omitted and these will reduce this discrepancy once properly propagated. Given the agreement between the PD and mock simulation analysis, there is nothing to suggest that the three independent trends highlighted in Colgain:2022rxy across OHD, Type Ia SNe and QSOs are not bona fide disagreements and that redshift evolution is present in the sample. The task remains to combine them at the level of a $\chi^{2}$ likelihood instead of combining them using Fisher’s method on the basis that they are independent probabilities. We leave this exercise for a forthcoming paper, but revisit the tension in OHD data in section 5. For completeness, in Table 2 we perform a reanalysis of CC data subsets with the PD approach and record the $68\%$ confidence intervals. In Fig. 6 we compare the $H_{0}$ constraints from Table 1 and Table 2, where it is clear that MCMC credible intervals shift to lower values at intermediate $z_{\textrm{min}}$.